A runner is practicing on a circular track that is 415 m in circumference. From the point farthest to the east on the track, he starts off running due south and follows the track as it curves around toward the west.
(a) If he runs halfway around the track and stops at the farthest western point of the track, what is the distance he traveled?
1
Expert's answer
2012-09-27T10:59:37-0400
The distance traveled is a half of circumference, or 207.5 m.
Otherwise, the
displacement will be equal to diameter D = C / Pi = 415/ 3.14159 = 132.09871
m.